1/2

Frédéric Bérat ,  Frédéric Bérat ,
Ma Normandie  My Normandy

translated by Tony Provencher

Quand tout renaît à l’espérance, Et que l’hiver fuit loin de nous, Sous le beau ciel de nôtre France, Quand le soleil revient plus doux, Quand la nature est reverdie, Quand l’hirondelle est de retour, J’aime à revoir ma Normandie! C’est le pays qui m’a donné le jour.

When far from us the Winter flies, When the world is born to hope anew, Under France’s lovely skies, When the sun returns in sweeter hue, When Nature ’round us greener be, When swallows homeward wing their way, I love to see my Normandy, The land that gave to me the light of day.

J’ai vu les champs de l’Helvétie, Et ses chalets et ses glaciers; J’ai vu le ciel de l’Italie, Et Venise et ses gondoliers. En saluant chaque patrie, Je me disais : aucun séjour N’est plus beau que ma Normandie! C’est le pays qui m’a donné le jour.

Switzerland’s dales I did behold, And her chalets and glaciers blue; I’ve seen Italian skies of gold, And the gondoliers of Venice too. Although these lands enchanted me, I knew there was no hideaway More lovely than my Normandy, The land that gave to me the light of day.

Il est un âge dans la vie, Où chaque rêve doit finir, Un âge où l’âme recueillie A besoin de se souvenir. Lorsque ma muse refroidie Aura fini ses chants d’amour, J’irai revoir ma Normandie! C’est le pays qui m’a donné le jour.

There comes a time, as life unfolds, That brings an end to reveries. A time when must the ransomed soul Revisit cherished memories. When time has chilled my muse and me, And songs of love are sung away, I’ll see again my Normandy, The land that gave to me the light of day.

…  …

2/2

Author of this French poem is Frédéric Bérat , published 1837.

Arg-1029-2086 (2014-05-29 14:20:40)

Tiu ĉi kanto-teksto troviĝas en http: // nuspel. org/ Normandie. html . Pri la verkinto vidu la vikipediejon http: // fr. wikipedia. org/ wiki/ FrГ�dГ�ric_BГ�rat . La kanto estas aŭdebla en http: // nuspel. org/ Normandie. html . Vidu ankaŭ: http: // manormandie. snoupi. com/ lyrics .

Translation of the French poem “Ma Normandie” by Frédéric Bérat into English by Tony Provencher.

Arg-1029-2088 (2014-05-29 19:58:50)

La kanto estas aŭdebla en http: // manormandie. snoupi. com/ music . The website of Mr. Provencher can be found at http: // tonysmusic. snoupi. com/ home2 .

http://www.poezio.net/version?poem-id=1029&version-id=2086,2088